# of our esteemed colleague Professor

### v1

@@ 1. Introduction
Given a smooth map $f\colon M\to N$ between manifolds of dimensions $m$ and $n$, respectively, with $m\geq n$, we say that $x\in M$ is a singular point if the differential $df_{x}\colon T_{x}M\to T_{f(x)}N$ does not have the maximal rank $n$. It is known that when $f$ is a generic smooth map into a surface $N=S$, the set $\Sigma(f)$ of its singular points consists of folds and cusps (defined in Section 2), see [16, 17]. More specifically, for a given closed manifold $M$ of $\dim{M}\geq 2$ and a surface $S$, the set of smooth maps $f\colon M\to S$ which have only folds and cusps as their singular points is open and dense in the space $C^{\infty}(M,S)$ of smooth maps of $M$ into $S$ endowed with the Whitney $C^{\infty}$ topology (for example, see [5]). In this sense, generic maps to surfaces can be considered as generalizations of Morse functions on manifolds which are generic maps into the $1$-dimensional manifold $\mathbb{R}$. Therefore, it is expected that the topology of the source manifold $M$ is somehow related to various properties of generic maps $M\to S$ on that manifold.
@@
This question is a generalization of the first of the two Takase problems formulated in [13, Problem 3.2]. Takase considered image simple generic maps $f\colon S^{3}\to S^{2}$ without definite fold singular points. (For the definition of a definite fold singularity, refer to Section 2.) The first Takase problem asks if the number $|\Sigma(f)|$ of components of $\Sigma(f)$ is congruent modulo two to $d+1$, where $d\in\mathbb{Z}\cong\pi_{3}(S^{2})$ is the homotopy class represented by $f$. The second problem asks if the minimum number of components of $\Sigma(f)$ is equal to $|d|+1$.
The second author [14] later answered the first Takase problem in the negative by constructing two homotopic image simple fold maps $f$ and $g$ without definite folds such that $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$, see Section 4 for details. Thus, in the case of image simple fold maps of manifolds of dimension $3$ to surfaces, the parity of the number of fold components is not a homotopy invariant. It is worth mentioning that the second author constructed an explicit homotopy $F\colon M\times[0,1]\to S\times[0,1]$ between $f$ and $g$. The singular point set of a generic homotopy between fold maps to a surface is, in general, a smooth surface properly embedded in $M\times[0,1]$, and we see that for the specific homotopy $F$ constructed in [14], the surface $\Sigma(F)$ of the singular points of $F$ is the disjoint union of $\Sigma(f)\times[0,1]$ and a Möbius band. Furthermore, the restriction of $F$ to the Möbius band has a triple self-intersection point.
@@
    the singular point set $\Sigma(F)$ of the homotopy is an orientable surface, or
@@
    the restriction $F|_{\Sigma(F)}$ of homotopy to its singular point set does not have triple self-intersection points, and either $M$ is a closed orientable manifold of dimension $3$ and $S$ is orientable, or $M$ is a closed manifold of odd dimension $m>2$ and $S$ is $\mathbb{R}^{2}$ or $S^{2}$.
@@
In fact, we will construct an explicit homotopy $F\colon M\times[0,1]\to S\times[0,1]$ satisfying the conclusion of Theorem 1.1. Its singular point set $\Sigma(F)$ contains a Möbius band component, and $F|_{\Sigma(F)}$ has a triple self-intersection point. This shows that conditions (2) and (3) above are necessary in general. On other other hand, since Theorem 1.1 permits the target surface $S$ to be non-orientable, it is not guaranteed that every homotopy satisfying the conclusion of the theorem must have a triple self-intersection point. However, if we remove the condition that the homotopy should be supported in a small disc, then, as we will see below, in the case where $M$ is even-dimensional, and both $M$ and $S$ are non-orientable, the absence of a triple self-intersection is not a necessary condition.
@@
We also construct an explicit homotopy $F$ between $f$ and $g$ in Theorem 1.2 such that the singular point set $\Sigma(F)$ is non-orientable, but that it has no triple self-intersections. This again shows that condition (2) above is necessary in general: however, condition (3) is not necessary in general.
To summarize, let $M$ be a closed manifold of dimension at least $2$, and $S$ be a surface. Let $F$ be a homotopy between two image simple fold maps $f$ and $g\colon M\to S$. Suppose that the dimension of $M$ is even.
@@
We do not know if the last statement remains true without the assumptions on $S$.
The paper is organized as follows. In Section 2 we prepare some basic notions necessary in the paper. In Section 3 we use open book decompositions as well as round fold maps of odd-dimensional spheres to prove Theorem 1.1 producing a homotopic modification $g$ of $f$ with $|\Sigma(g)|\equiv\#|\Sigma(f)|+1\pmod{2}$. In Section 4 we give another proof of Theorem 1.1 along the argument in [14], using allowable moves (or always-realizable moves) as listed in [15]. In Section 5, we first give a new proof of the homotopy invariance of the parity of the number of connected components of the singular point set for image simple fold maps from a closed even-dimensional manifold to an orientable surface (see Theorem 5.1). Then, we prove Theorem 1.2 by constructing an explicit example.
@@ 3. A peculiar open book decomposition of the sphere $S^{2n+1}$
For $v\in T_{x}S^{n}$ with $v||\leq\pi$ and $t\in[0,1]$, let $\gamma_{t}(x,v)$ be the point $(\ell(t),\ell^{\prime}(t))\in P$, where $\ell$ is the unique geodesic as described above. We see easily that this defines a well-defined smooth map $\gamma_{t}\colon P\to P$. By an argument similar to the above, we see that this defines a smooth $1$-parameter family of diffeomorphisms. Furthermore, we see that $\gamma_{0}$ is the identify of $P$ and that $\gamma_{1}=\gamma$. This completes the proof. ∎
@@
Now, let $f\colon M\to S$ be an arbitrary image simple fold map into a (possibly non-orientable) surface $S$. We first consider the disjoint union $f\cup f_{X}\colon M\cup X\to S$, where $f_{X}(X)$ is contained in a small $2$-disc in $f(M)\setminus f(\Sigma(f))$. We apply the birth move (see [15, Fig. 4]) to create a wrinkle consisting of two cusps together with fold arcs of absolute indices 0 and $1$ in such a way that the image sits just next to $f_{X}(\Sigma(f_{X}))$. (For this and the following, see Fig. 3. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows.) Then, we construct an image simple fold map $M\#X\cong M\to S$ by the connected sum operation along definite folds, as described in [12, Lemma 5.4], in such a way that the image of the singular point set is as depicted in the upper right figure of Fig. 3. Note that the fold index $n$ does not change even if we change the normal direction, since the dimension of the source manifold is equal to $2n+1$. We apply the move of type II twice so that the image of the fold circle of index $n$ gets out of the wrinkle. This is possible as long as n<2n1n<2n-1, i.e. $n\geq 2$ (see [15, Fig. 5]). Finally, we can eliminate the wrinkle, which is possible, since one of the fold arcs has absolute index 0 (see [4]). The resulting fold map $g\colon M\to S$ is image simple and we have $|\Sigma(g)|=\#|\Sigma(f)|+1$.
@@
(2) We see easily that $f$ and $g$ are homotopic as continuous maps. However, we do not know how to construct explicit homotopy between $f$ and $g$ thus constructed.
@@ 4. Allowable moves approach
Let us give another proof to Theorem 1.1 by an approach similar to that in [14]. For the moves that we use in this section, the reader is referred to [15].
Suppose $n\geq 2$. Let $f\colon M\to S$ be an arbitrary image simple fold map of a closed $(2n+1)$-dimensional manifold into a (possibly non-orientable) surface $S$. We consider a small $2$-disc $\Delta$ contained in $f(M)\setminus f(\Sigma(f))$, and we modify the map $f$ on a connected component of $f^{-1}(\Delta)$, following the procedure as described in Fig. 4. These figures depict the singular value sets of generic maps inside $\Delta$. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows. The lower left figure depicts the singular value set inside the $2$-disc $\Delta$, i.e. the empty set. Then, we create three wrinkles, using the birth move three times (see [15, Fig. 4]). Then, by three flip moves, we turn the wrinkles into three bow ties with involved folds of indices n1n-1, $n$ and $n+1$ (see [15, Fig. 4]). Note that the index n1n-1 turns into $n+1$ if we reverse the normal direction. Then, by three cusp merges we merge the three bow ties as in the upper right figure. This is possible, since we are modifying the maps only on a connected component of $f^{-1}(\Delta)$ so that the relevant fibers over the points on the paths between the images of the cusps that are merged are connected. As the three fold curves all have indices $n\geq 2$, we can perform the move of type III as in the lower right figure (see [15, Fig. 6]). Finally, we need to perform three unflips. This is possible, since the involved folds are of indices $n$ and $n+1$ with $2\leq n\leq 2n-1$ (see [15, Fig. 4]).
[…]
@@
The above argument may not work directly for $n=1$. However, as noted in [14], by merging the three pairs of cusps with appropriate “twists”, we can guarantee that the move of type III can be applied.
@@
This approach allows us to determine the singular point set of the constructed homotopy as follows. As we perform three births, we have three $2$-discs as the singular point set in the first three steps. In the fourth step, we merge the $2$-discs by three bands in such a way that the resulting compact surface has exactly one boundary circle whose image is depicted in the upper right figure of Fig. 4. Therefore, the singular point set for the first four steps must be a Möbius band. The rest of the homotopy does not change the topological type of the singular point set component. Therefore, the resulting singular point set of the homotopy is a Möbius band, except for the cylinder part corresponding to the fold circles of the original image simple fold map $f$.
So, the singular point set of the entire homotopy $F\colon M\times[0,1]\to S\times[0,1]$ is diffeomorphic to the disjoint union of $|\Sigma(f)|$ copies of the cylinder together with a Möbius band. Note also that $F|_{\Sigma(F)}$ has a triple self-intersection point.
@@
Let us give a short remark concerning the first Takase problem formulated in [13, Problem 3.2] mentioned in Section 1. The map $f\colon S^{3}\to\mathbb{R}^{2}$ constructed in [14] does not give a direct negative answer to the problem. However, if we use the procedure as described above, we get a counter example for every element $n$ of $\pi_{3}(S^{2})\cong\mathbb{Z}$, since the newly created fold circle is indefinite. Therefore, the number modulo two of fold components of such image simple fold maps without definite fold points is not a homotopy invariant in general.

### v2

# This paper is dedicated to the memory of our esteemed colleague Professor

@@ 1. Introduction
Given a smooth map $f\colon M\to N$ between manifolds of dimensions $m$ and $n$, respectively, with $m\geq n$, we say that $x\in M$ is a singular point if the differential $df_{x}\colon T_{x}M\to T_{f(x)}N$ does not have the maximal rank $n$. It is known that when $f$ is a generic smooth map into a surface $N=S$, the set $\Sigma(f)$ of its singular points consists of folds and cusps (defined in Section 2), see [21, 22]. More specifically, for a given closed manifold $M$ of dimension $\geq 2$ and a surface $S$, the set of smooth maps $f\colon M\to S$ which have only folds and cusps as their singular points is open and dense in the space $C^{\infty}(M,S)$ of smooth maps of $M$ into $S$ endowed with the Whitney $C^{\infty}$ topology (for example, see [7]). In this sense, generic maps to surfaces can be considered as generalizations of Morse functions on manifolds which are generic maps into the $1$-dimensional manifold $\mathbb{R}$. Therefore, it is expected that the topology of the source manifold $M$ is somehow related to various properties of generic maps $M\to S$ on that manifold.
@@
Example 1.1.
The restriction of the standard projection $\mathbb{R}^{3}\to\mathbb{R}^{2}$ to the standard unit sphere defines an image simple fold map $f\colon S^{2}\to\mathbb{R}^{2}$ with $|\Sigma(f)|=1$. For any prescribed positive odd integer $k$, it is easy to construct an image simple fold map $f_{k}$ of $S^{2}$ into $\mathbb{R}^{2}$ with $|\Sigma(f_{k})|=k$; however, no such map exists for any positive even integer $k$ by [10, Theorem 1.1] or by Theorem 6.1 of the present paper. In fact, according to [23], there exists no fold map $g\colon S^{2}\to\mathbb{R}^{2}$ with $|\Sigma(g)|$ even, even if we allow $g$ not to be image simple. See also Remark 6.3.
The above question is a generalization of the first of the two Takase problems formulated in [18, Problem 3.2]. Takase considered image simple generic maps $f\colon S^{3}\to S^{2}$ without definite fold singular points with $\Sigma(f)\neq\emptyset$. (For the definition of a definite fold singularity, refer to Section 2.) The first Takase problem asks if the number $|\Sigma(f)|$ of components of $\Sigma(f)$ is congruent modulo two to $d+1$, where $d\in\mathbb{Z}\cong\pi_{3}(S^{2})$ is the homotopy class represented by $f$. The second problem asks if the minimum number of components of $\Sigma(f)$, over all $f$ as above in the given homotopy class $d\in\mathbb{Z}\cong\pi_{3}(S^{2})$, is equal to $|d|+1$.
The second author [19] later answered the first Takase problem in the negative by constructing two homotopic image simple fold maps $f$ and $g$ without definite folds such that $|\Sigma(f)|\not\equiv\#|\Sigma(g)|\pmod{2}$, see Section 5 for details. Thus, in the case of image simple fold maps of manifolds of dimension $3$ to surfaces, the parity of the number of fold components is not a homotopy invariant. It is worth mentioning that the second author constructed an explicit generic homotopy $F\colon M\times[0,1]\to S\times[0,1]$ between $f$ and $g$. The singular point set of a generic homotopy between fold maps to a surface is, in general, a smooth surface properly embedded in $M\times[0,1]$, and we see that for the specific generic homotopy $F$ constructed in [19], the surface $\Sigma(F)$ of the singular points of $F$ is the disjoint union of $\Sigma(f)\times[0,1]$ and a Möbius band. Furthermore, the restriction of $F$ to the Möbius band has a triple self-intersection point.
Note that every homotopy between image simple fold maps can be approximated arbitrarily well by a generic homotopy (for details, see §3). Therefore, when two image simple fold maps are homotopic, we may assume that the homotopy is generic.
@@
    $F$ is generic and the singular point set $\Sigma(F)$ of the homotopy is an orientable surface, or
@@
    $F$ is generic and the restriction $F|_{\Sigma(F)}$ of homotopy to its singular point set does not have triple self-intersection points, and either $M$ is a closed orientable manifold of dimension $3$ and $S$ is orientable, or $M$ is a closed manifold of odd dimension $m>2$ and $S$ is $\mathbb{R}^{2}$ or $S^{2}$.
@@
In fact, we will construct an explicit generic homotopy $F\colon M\times[0,1]\to S\times[0,1]$ satisfying the conclusion of Theorem 1.2. Its singular point set $\Sigma(F)$ contains a Möbius band component, and $F|_{\Sigma(F)}$ has a triple self-intersection point. This shows that conditions (2) and (3) above are necessary in general. On other hand, since Theorem 1.2 permits the target surface $S$ to be non-orientable, it is not guaranteed that every generic homotopy satisfying the conclusion of the theorem must have a triple self-intersection point. However, if we remove the condition that the homotopy should be supported in a small disc, then, as we will see below, in the case where $M$ is even-dimensional, and both $M$ and $S$ are non-orientable, the absence of a triple self-intersection is not a necessary condition.
@@
We also construct an explicit generic homotopy $F$ between $f$ and $g$ in Theorem 1.3 such that the singular point set $\Sigma(F)$ is non-orientable, but that it has no triple self-intersections. This again shows that condition (2) above is necessary in general: however, condition (3) is not necessary in general.
To summarize, let $M$ be a closed manifold of dimension at least $2$, and $S$ be a surface. Let $F$ be a (generic) homotopy between two image simple fold maps $f$ and $g\colon M\to S$. Suppose that the dimension of $M$ is even.
@@
Note that the last statement does not hold in general without the assumption on $S$ according to Theorem 1.3.
The paper is organized as follows. In Section 2 we prepare some basic notions necessary in the paper. In Section 3, for the reader’s convenience, based on [5, 8, 10, 20], we summarize the moves for generic maps together with sufficient conditions for some of them to be allowable, which will be used in later sections for constructing explicit homotopies. In Section 4 we use open book decompositions as well as round fold maps of odd-dimensional spheres to prove Theorem 1.2 producing a homotopic modification $g$ of $f$ with $|\Sigma(g)|\equiv\#|\Sigma(f)|+1\pmod{2}$. In Section 5 we give another proof of Theorem 1.2 along the argument in [19], using allowable moves (or always-realizable moves) as listed in Section 3. In Section 6, we first give a new proof of the homotopy invariance of the parity of the number of connected components of the singular point set for image simple fold maps from a closed even-dimensional manifold to an orientable surface (see Theorem 6.1). Then, we prove Theorem 1.3 by constructing an explicit example.
@@ 2. Preliminaries
3. Generic homotopies of maps to a surface
Let $C^{\infty}(M,N)$ denote the space of smooth maps of a manifold $M$ into a manifold $N$. It is endowed with the Whitney $C^{\infty}$ topology. We say that a map $f\colon M\to N$ is right-left equivalent to a map $g\colon M\to N$ if $g=\Phi\circ f\circ\Psi^{-1}$ for some diffeomorphisms $\Psi$ of $M$ and $\Phi$ of $N$. The same definition applies to map germs. A map $f\colon M\to N$ is said to be stable if every map $g$ sufficiently close to $f$ in $C^{\infty}$ topology is right-left equivalent to $f$. A map germ of $f\colon M\to N$ is stable at a point $p\in M$ if for every neighborhood $U$ of $p$ and every map $g\colon M\to N$ sufficiently close to $f$ there is a point $p^{\prime}$ in $U$ such that the map germ of $g$ at $p^{\prime}$ is right-left equivalent to the map germ of $f$ at $p$ (for example, see [1]).
Let $f\colon M\to N$ be a smooth map. For a point $p$ in $M$, we define $\Sigma_{p}(f)$ as the set of points in $M$ at which the map germ of $f$ is right-left equivalent to the map germ of $f$ at $p$. Mather [14] proved that if all map germs of $f$ are stable, then the sets $\Sigma_{p}(f)$ are submanifolds of $M$. For a finite family of subspaces $\{P_{i}\}_{i=1}^{s}$ of a vector space, we say that the spaces are in general position if for every sequence of integers $1\leq i_{1}<i_{2}<\cdots<i_{r}\leq s$, we have
$\mathrm{codim}\,(P_{i_{1}}\cap P_{i_{2}}\cap\cdots\cap P_{i_{r}})=\sum_{k=1}^{r}\mathrm{codim}\,P_{i_{k}},$
where $\mathrm{codim}$” denotes the codimension. Suppose that all map germs of $f$ are stable and for any finitely many points $p_{1},p_{2},\ldots,p_{s}$ with $s>1$ and $f(p_{1})=f(p_{2})=\cdots=f(p_{s})=q$, the subspaces $d_{p_{i}}f(T_{p_{i}}\Sigma_{p_{i}}(f))$ are in general position in $T_{q}N$. Then we say that $f$ satisfies the Mather normal crossing condition.
In a series of papers, Mather investigated stability of maps and proved its equivalence to various other conditions. In particular, Mather proved [14] that a proper map $f$ is stable if and only if it satisfies the Mather normal crossing condition (for this particular interpretation of the Mather result, and a new proof, see [16]). In the case of maps with fold and cusp singularities the Mather normal crossing condition simply requires that the images of submanifolds of fold and cusp singular points intersect (and self-intersect) generically. In particular, image simple fold maps are always stable.
To study homotopies of stable maps, we work with generic homotopies. For the definition and some properties of generic homotopies, the reader is referred to [10]. In particular, the class of generic homotopies is open and dense in the space of all homotopies (e.g., see [10, Lemma 4.1]), which means that an arbitrary homotopy between stable maps can be approximated arbitrarily well by a generic homotopy. For this reason, we may and will almost always assume that a homotopy between stable maps is generic. Furthermore, under a generic homotopy, the image of the singular point set is modified by isotopy as well as by finitely many local modifications or moves listed in Theorem 3.1 below (e.g., see [10, Theorem 4.2]). For a precise formulation of moves and their realizability (or always-realizability), the reader is referred to [20, §3] as well.
Theorem 3.1.
Under a generic homotopy between stable maps of a closed manifold of dimension $\geq 2$ to a surface, the image of the singular point set is modified by isotopy as well as
  • type II fold crossings and their reverse moves,
  • type III fold crossings,
  • cusp-fold crossings and their reverse moves,
  • birth-death moves,
  • cusp merges and fold merges, and
  • flip-unflip moves.
To discuss the realizability of the moves listed in Theorem 3.1, we need a coorientation (or normal orientation) convention as follows. Suppose that $p$ is a fold singular point of $f$, $(t,x_{1},x_{2},\ldots,x_{m-1})$ are local coordinates around $p$, and $(X,Y)$ are local coordinates around $f(p)$ with respect to which $f$ is given by its normal form as in Definition 2.1. We choose the coorientation of $f(\Sigma(f))$ at $f(p)$ as indicated in Fig. 2. We emphasize that locally in the standard coordinate chart, the coorientation of the image of the fold curve is chosen in the direction of Y-Y.
Figure 2. Coorientation convention at a fold singularity
Now we discuss the six local modifications listed in Theorem 3.1.
Type II fold crossing
The type II fold crossing increases the number of self-intersection points of $f(\Sigma(f))$ by two, see Fig. 3 from left to right. Suppose that the two fold curves are of indices $i$ and $j$ as indicated in the figure. If $i\leq j$, then the fold crossing move is always-realizable by a (generic) homotopy. The inverse to the fold crossing is realizable if $j<i$.
Figure 3. Type II fold crossing and its reverse move
Type III fold crossing
This move occurs at a triple self-intersection point of the images of three arcs of fold singular points. Suppose that the three fold curves are of indices $i,j$ and $k$ as in the left of Fig. 4. The move in the figure from left to right is realizable if any of the following three conditions are satisfied (see [5, Chapter IV, §2.2, Prop. 2]):
  • $i+k\geq m$,
  • $\max\{i,k\}>j$, or
  • $i_{1}=i_{2}=i_{3}\geq 2$.
The move from right to left is realizable if
  • $i+k\leq m-2$,
  • $\min\{i,k\}<j$, or
  • $i=j=k\leq m-3$.
Figure 4. Type III fold crossing
Cusp-fold crossing
The cusp-fold crossing is shown in Fig. 5, from left to right. This move is realizable if $i>0$. The reverse move is realizable if (see [5, Chapter IV, §3.3, Prop. 4])
  • i+j<m2i+j<m-2, or
  • $i<j$.
Figure 5. Cusp-fold crossing and its reverse move
Birth move and death move
The birth move creates a new component of singular points. Such a move is always-realizable, see Fig. 6. Its reverse, i.e., the move that eliminates a component is always-realizable in the case where one of the fold arcs is of absolute index 0, see [5].
Figure 6. Birth move (from left to right) and death move (from right to left)
Cusp merge and fold merge
The cusp merge, depicted in Fig. 7, from left to right, is realizable provided that the fiber over the point $\ast$ is path connected. The fold merge, which is depicted in Fig. 7, from right to left, is not always-realizable.
Figure 7. Cusp merge (from left to right) and fold merge (from right to left)
Flip and unflip
The flip move is depicted in Fig. 8, from left to right. It is always-realizable for i<m1i<m-1. The unflip is depicted in Fig. 8, from right to left. It is realizable if $i=0$ or $2\leq i\leq m-2$, see [8, Part I, Chapter V, Proposition 1.4].
Figure 8. Flip (from left to right) and unflip (from right to left)
@@ 4. A peculiar open book decomposition of the sphere $S^{2n+1}$
Example 4.1.
In the case $n=1$, the disc bundle $P$ over $S^{1}$ is diffeomorphic to the cylinder $S^{1}\times[0,2\pi]$. We will define a generalized Dehn twist motivated by the standard example. The standard Dehn twist is a diffeomorphism of the cylinder that twists it fixing the boundary pointwise. In coordinates it is given by $(x,t)\mapsto(xe^{it},t)$.
@@
For $v\in T_{x}S^{n}$ with $v||\leq\pi$ and $t\in[0,1]$, let $\gamma_{t}(x,v)$ be the point $(\ell(t),\ell^{\prime}(t))\in P$, where $\ell$ is the unique geodesic as described above. We see easily that this defines a well-defined smooth geodesic flow $\gamma_{t}\colon P\to P$, which will be often considered as an isotopy of $P$, in the following argument. By an argument similar to the above, we see that this defines a smooth $1$-parameter family of diffeomorphisms. Furthermore, we see that $\gamma_{0}$ is the identify of $P$ and that $\gamma_{1}=\gamma$. This completes the proof. ∎
@@
Now, let $f\colon M\to S$ be an arbitrary image simple fold map into a (possibly non-orientable) surface $S$. We first consider the disjoint union $f\cup f_{X}\colon M\cup X\to S$, where $f_{X}(X)$ is contained in a small $2$-disc in $f(M)\setminus f(\Sigma(f))$. We apply the birth move (see Fig. 6) to create a wrinkle consisting of two cusps together with fold arcs of absolute indices 0 and $1$ in such a way that the image sits just next to $f_{X}(\Sigma(f_{X}))$. (For this and the following, see Fig. 10. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows, see Section 3. Then, we construct an image simple fold map $M\#X\cong M\to S$ by the connected sum operation along definite folds, as described in [17, Lemma 5.4], in such a way that the image of the singular point set is as depicted in the upper right figure of Fig. 10. Note that the fold index $n$ does not change even if we change the normal direction, since the dimension of the source manifold is equal to $2n+1$. We apply the reverse move of type II crossing twice so that the image of the fold circle of index $n$ gets out of the wrinkle. This is possible as long as n<2n1n<2n-1, i.e. $n\geq 2$ (see Fig. 3). Finally, we can eliminate the wrinkle, which is possible, since one of the fold arcs has absolute index 0 (see Fig. 6).
The resulting fold map $g\colon M\to S$ is image simple and we have $|\Sigma(g)|=\#|\Sigma(f)|+1$.
@@
(2) We see easily that $f$ and $g$ are homotopic as continuous maps. However, we do not know how to construct an explicit homotopy between $f$ and $g$ thus constructed.
@@ 5. Allowable moves approach
Let us give another proof to Theorem 1.2 by an approach similar to that in [19]. For the moves that we use in this section, the reader is referred to Section 3.
Suppose $n\geq 2$. Let $f\colon M\to S$ be an arbitrary image simple fold map of a closed $(2n+1)$-dimensional manifold into a (possibly non-orientable) surface $S$. We consider a small $2$-disc $\Delta$ contained in $f(M)\setminus f(\Sigma(f))$, and we modify the map $f$ on a connected component of $f^{-1}(\Delta)$, following the procedure as described in Fig. 11. These figures depict the singular value sets of generic maps inside $\Delta$. The integers attached to some curves indicate the corresponding indices of the relevant folds with respect to the normal directions indicated by arrows. The lower left figure depicts the singular value set inside the $2$-disc $\Delta$, i.e. the empty set. Then, we create three wrinkles, using the birth move three times (see Fig. 6). Then, by three flip moves, we turn the wrinkles into three bow ties with involved folds of indices n1n-1, $n$ and $n+1$ (see Fig. 8). Note that the index n1n-1 turns into $n+1$ if we reverse the normal direction. Then, by three cusp merges we merge the three bow ties as in the upper right figure (see Fig. 7). This is possible, since we are modifying the maps only on a connected component of $f^{-1}(\Delta)$ so that the relevant fibers over the points on the paths between the images of the cusps that are merged are connected. As the three fold curves all have indices $n\geq 2$, we can perform the type III crossing as in the lower right figure (see Fig. 4). Finally, we need to perform three unflips. This is possible, since the involved folds are of indices $n$ and $n+1$ with $2\leq n\leq 2n-1$ (see Fig. 8).
@@
The above argument may not work directly for $n=1$. However, as noted in [19], by merging the three pairs of cusps with appropriate “twists”, we can guarantee that the type III crossing can be applied.
@@
This approach allows us to determine the singular point set of the constructed homotopy as follows. As we perform three births, we have three $2$-discs as the singular point set in the first three steps. In the fourth step, we merge the $2$-discs by three bands in such a way that the resulting compact surface has exactly one boundary circle whose image is depicted in the upper right figure of Fig. 11. Therefore, the singular point set for the first four steps must be a Möbius band. The rest of the homotopy does not change the topological type of the singular point set component. Therefore, the resulting singular point set of the (generic) homotopy is a Möbius band, except for the cylinder part corresponding to the fold circles of the original image simple fold map $f$.
So, the singular point set of the entire (generic) homotopy $F\colon M\times[0,1]\to S\times[0,1]$ is diffeomorphic to the disjoint union of $|\Sigma(f)|$ copies of the cylinder together with a Möbius band. Note also that $F|_{\Sigma(F)}$ has a triple self-intersection point.
@@
Let us give a short remark concerning the first Takase problem formulated in [18, Problem 3.2] mentioned in Section 1. The map $f\colon S^{3}\to\mathbb{R}^{2}$ constructed in [19] does not give a direct negative answer to the problem. However, if we use the procedure as described above, we get a counter example for every element $n$ of $\pi_{3}(S^{2})\cong\mathbb{Z}$, since the newly created fold circle is indefinite. Therefore, the number modulo two of fold components of such image simple fold maps without definite fold points is not a homotopy invariant in general.